A rock group is going on a talk show. Their Song is 5 minutes 22 seconds, but their time slot is only 4 minutes 58 seconds long. How much faster do they have to play to get the song to fit their time slot?

Answer by nerdybill(7384)

A rock group is going on a talk show. Their Song is 5 minutes 22 seconds, but their time slot is only 4 minutes 58 seconds long. How much faster do they have to play to get the song to fit their time slot?
.
First, convert time in seconds:
5 minutes 22 seconds = (5*60)+22 = 300+22 = 322 seconds
4 minutes 58 seconds = (4*60)+58 = 240+58 = 298 seconds
.
What percent of time is available?
298/322 * 100 = 0.93 * 100 = 93%
This says that only 93% of the time is available.
.
Therefore, the group has to play:
100-93 = 7% faster